# The Mechanics of Nucleation and Growth and the Surface Tensions of Active Matter

### v1

@@ II.2 Mechanics of a Critical Nucleus
Given a saturation density, one may use Eqs. (6) to solve for the unknowns $\rho^{\rm in}$ and $R_{C}$. We note that as $\rho^{\rm sat}$ approaches either binodal density, $R_{C}$ will diverge; Eq. (4a) and Eq. (5a) then simplify to the macroscopic phase coexistence criteria presented in Ref. [35]. As one approaches this limit and $R_{C}$ grows, $\mathcal{O}\left(R_{C}^{-2}\right)$ corrections to the coexistence criteria become negligible and we expect our theory to be most accurate. Solution of the coexistence criteria is complicated by the presence of $\upgamma_{\rm mech}$ and $\upgamma_{\rm ost}$ in Eqs. (6), which depend on the density profile, which in turn depends on $R_{C}$ and $\rho^{\rm in}$. Therefore, solution of Eqs. (6) requires self-consistently solving for the density profile using a procedure detailed in Section 1 of the SI [33]. In the following section, we apply our mechanical theory to a canonical model system for active matter, active Brownian particles.
@@ II.3 Nucleation of Motility-Induced Phases
The form of $\mathcal{P}(\rho)$ and $\{\kappa_{i}(\rho)\}$ for ABPs are given by [35, 26]:
$\mathcal{P}(\rho)$ $=p_{C}(\rho)+\frac{\rho\zeta U_{o}\ell_{o}\overline{U}(\rho)}{d(d-1)},$ (8a)
$\kappa_{1}(\rho)$ $=\frac{3\ell_{o}^{2}}{2d(d-1)(d+2)}\left(\overline{U}(\rho)\right)^{2}\frac{\partial p_{C}}{\partial\rho},$ (8b)
$\kappa_{3}(\rho)$ $=\frac{3\ell_{o}^{2}}{2d(d-1)(d+2)}\overline{U}(\rho)\frac{\partial}{\partial\rho}\left[\overline{U}(\rho)\frac{\partial p_{C}}{\partial\rho}\right],$ (8c)
@@
$\mathcal{K}_{r}$ $=\mathcal{K}_{t}=0,$ (8e)
where $p_{C}(\rho)$ is the conservative pressure arising from particle interactions, $\overline{U}(\rho)$ is the dimensionless effective active speed, and $d(>1)$ is the system dimensionality. In Eqs. (8) we have ignored the gradient contributions of the interaction potential to $\bm{\Sigma}$, which is justified when $\ell_{o}/d_{\rm hs}>>1$ for purely repulsive interactions [30, 35, 44]. Given our expressions for $\{\kappa_{i}(\rho)\}$, one can solve for the generalized Maxwell construction pseudovariable $\mathcal{E}\sim p_{C}(\rho)$ [35]. We also note that because $\mathcal{K}_{r}=0$, Eq. (4a) reduces to the familiar Laplace pressure difference (i.e. no non-local corrections) and the negative sign of $\upgamma_{\rm mech}$ [21] implies that the pressure inside of a droplet is lower than the pressure outside of the droplet [30]. The phase diagram, dependent on $\ell_{o}/d_{\rm hs}$ and volume fraction $\phi\equiv\pi d_{\rm hs}^{3}/6$, has been well established computationally [5] and theoretically [35] with the latter predicting a critical point at $\phi\approx 0.49$ and $\ell_{o}\approx 16.2d_{\rm hs}$
Figure 2: Critical radius and relative change in nucleus density (${\nu^{\rm in}\equiv|\rho^{\rm in}-\rho^{\rm gas,\rm liq}|/\rho^{\rm gas,\rm liq}}$) as a function of supersaturation (${\nu^{\rm sat}\equiv|\rho^{\rm sat}-\rho^{\rm gas,\rm liq}|/\rho^{\rm gas,\rm liq}}$).
@@
Figure 2 displays a striking difference between critical bubbles and droplets. At all examined values of activity, critical droplets of active Brownian particles are over an order of magnitude larger than critical bubbles under similar saturation conditions. This difference between bubbles and droplets widens as activity is increased and the phases become increasingly distinct. To contextualize this order of magnitude difference, we calculate the critical nucleus size and density for an equilibrium van der Waals fluid [33]. While the size of the critical droplets are indeed larger than the size of critical bubbles in passive fluids, they remain within an order of magnitude of each other at similar saturation conditions even far from the critical point. In equilibrium, this difference is rooted in the asymmetry of the equations of state outside of the binodal densities: the vapor bubbles are more compressible than the liquid droplets. This asymmetry results in larger Laplace pressure differences for bubbles in comparison to droplets under similar saturation conditions, leading to the critical radius discrepancy. While this asymmetry in the compressibility is also present for active systems, the coexistence criteria contain uniquely nonequilibrium contributions, including a significant discrepancy between the vapor and bubble Ostwald tensions. At equilibrium these two quantities are identical, but can differ by an order of magnitude in active systems as shown in Fig. 3.
@@ II.4 Langevin Dynamics of a Nucleus
where $\rho$ now is a stochastic density field (while in the previous sections it was an ensemble average), $\bm{\Sigma}$ is defined identically as the dynamic stress tensor of the previous section and $\bm{\eta^{\rm act}}$ is an athermal stochastic contribution to the flux with zero mean and a variance of:
@@
We recently established a surface-area minimizing principle for liquid-gas interfaces of ABPs resulting from a positive capillary-wave tension, $\upgamma_{\rm cw}$, which resists fluctuations to the area-minimizing interfacial shape. The strength of the fluctuations were found to be set by the athermal active energy scale $k_{B}T^{\rm act}$ [26]. The ratio of the capillary tension to this energy scale – the active interfacial stiffness – controls the amplitude of long-wavelength interfacial fluctuations. A fluctuating nucleus can be well-approximated as a sphere when the average radius $R$ satisfies $(R^{2}\upgamma_{\rm cw}/k_{B}T^{\rm act})^{1/2}>>1$. Under this condition, the stochastic density field dynamics can be connected to the radial growth dynamics of the nucleus through the following radially symmetric ansatz [46, 4]:
@@
The radial growth dynamics take the form of a one-dimensional overdamped Langevin equation containing three distinct forces: a drag force with drag coefficient $\zeta_{\rm eff}$, a stochastic force with an athermal temperature scale of $k_{B}T^{\rm act}$ [47], and a deterministic force $F^{D}(R)$ proportional to the Ostwald tension. This deterministic force vanishes when $R=R_{C}$ where the density outside nucleus is identical to the far away saturation density $\rho^{\rm out}=\rho^{\rm sat}$. Crucially, the noise statistics satisfy the fluctuation-dissipation theorem (FDT) despite the underlying microscopic dynamics [Eqs. (7)] violating FDT. We can plainly see that $\upgamma_{\rm ost}$ is the quantity controlling the dynamics of nucleation and ripening. As first argued by Thjung et al. [23] a positive $\upgamma_{\rm ost}$ implies $R_{C}$ is an unstable fixed point in which nuclei with $R>R_{C}$ tend to grow while nuclei with $R<R_{C}$ tend to shrink. A priori, there is no reason to assume $\upgamma_{\rm ost}$ is generally positive. In the case of a negative $\upgamma_{\rm ost}$, $R_{C}$ becomes a stable fixed point. However, for ABPs we do not find a negative $\upgamma_{\rm ost}$ [see Fig. 4] – implying that nucleation occurs with the same qualitative dynamics as an equilibrium system as observed in simulation [2].
@@
with $F^{D}(R)=-\partial U/\partial R$. In defining this potential, we have treated $\upgamma_{\rm ost}$ as independent of $R$ (just as in Ref. [4]), an approximation that appears to be well justified while $R\sim R_{C}$ [33]. This potential may be rearranged into a form familiar to equilibrium intuition [33]:
@@
At equilibrium, $\mathcal{V}$ simplifies to the expected $\mu(\rho^{\rm sat})\Delta\rho-\Delta f$ [33]. From Eq. (13), one can clearly appreciate that the barrier contains two contributions: a positive term which depends solely on interfacial properties scaling as $R^{2}$ and a negative term which depends solely on bulk properties scaling as $R^{3}$. Thus the familiar picture of homogeneous nucleation where interfacial effects compete with bulk effects in order to define a critical radius and nucleation barrier is recovered but now contains generalized volumetric driving and interfacial penalty terms.
@@
The recovery of an effective detailed balance with respect to the size of a sub-critical nucleus directly implies that the probability of observing a nucleus of size $R<R_{C}$ is proportional to ${\text{exp}\left[-U(R)/k_{B}T^{\rm act}\right]}$ [4]. We note that the potential defined here relied on the assumption that the density profiles and quantities such as $\upgamma_{\rm ost}$ are constant with $R$ an approximation that clearly breaks down when $R<<R_{C}$. Even in equilibrium, the nucleation potential formulated using bulk thermodynamic quantities (e.g  the so-called “capillary approximation”) is often used for all $R<R_{C}$ despite likely being beyond the range of applicability of bulk thermodynamic quantities [12, 48, 49]. Nevertheless, use of this potential for all $R$ allows one to straightforwardly extend the arguments of Becker-Döring theory [9, 12, 15] to extract the following kinetic expression for the rate at which an active droplet nucleates per unit volume [33]:
@@
The characteristic timescale to observe a nucleation event is given by $\tau_{\rm nuc}=(V\mathcal{R})^{-1}$ where $V$ is the total volume of the system. Thus the work of Redner et al. [2], which inferred the existence of a nucleation barrier by observing clustering statistics consistent with Becker-Döring theory from particle simulation, is well-justified on theoretical grounds. We note that the rate [Eq. (14)] is applicable to the nucleation of active droplets but not to active bubbles – the physical interpretation of the “clusters” posited in the derivation of Eq. (14) becomes unclear [12] in this context.
@@ II.5 Active Coarsening Dynamics
While Eqs. (6) and Eqs. (11) were motivated to describe the earliest stages of nucleation, i.e. the structure and dynamics of a single nucleus forming in a supersaturated solution, they may also describe late-stage coarsening dynamics long after all nucleation events have occurred. At this stage, the system will contain several nuclei of different sizes. The coarsening of these domains can proceed through a combination of several different mechanisms, including evaporation-condensation (Ostwald ripening) and coalescence. At sufficiently low $\nu^{\rm sat}$, the nuclei will be separated by large enough distances that collisions between them are exceedingly rare and coarsening will be dominated by Ostwald ripening. In such a scenario, Eqs. (11) can describe the dynamics of each nucleus such that they are only coupled via the density of the majority phase, $\rho^{\rm sat}$, which now evolves with time. When $\upgamma_{\rm ost}>0$, we expect the forward Ostwald ripening process to occur and larger nuclei of minority phase to grow at the expense of smaller nuclei until the system reaches a final steady state with a single macroscopic domain. Nuclei compete to absorb (or inject) particles from the majority phase and therefore $R_{C}$, the radius below which nuclei shrink and above which nuclei grow, is now a function of time.
With the above physical picture in mind, we can determine the time-dependence of the post-nucleation critical radius in a similar manner as the Lifshitz-Slyozov-Wagner (LSW) theory [11]. In the limit of a vanishing phase fraction of the minority phase, the nuclei can be approximated as independent with each satisfying the Langevin equation shown in Eq. (11). The Fokker-Planck equation consistent with Eq. (11) can be straightforwardly found and its long-time asymptotic solution (while discarding terms of higher order than $R^{-3}$) can be determined analogously to LSW theory, as detailed in Section 4 of the SI [33]. In doing so, we find that the critical radius evolves in time according to:
@@
where $\mathcal{C}$ is a dimensionless constant and $t_{o}$ is an arbitrary time origin. We note that this long-time asymptotic solution also implies that the average radius is always within a dimensionless multiplicative constant of $R_{C}$ [33, 11, 14]. Equation (15) demonstrates that the $t^{1/3}$ growth law predicted by LSW theory should generically hold out of equilibrium while $\upgamma_{\rm ost}>0$. This $1/3$ power law has been measured experimentally in active systems [6, 7] with disparate microscopic details, suggesting that this growth law may in fact be robust. We note that the simulations of Redner et al. [8] measured the average number of particles per droplet (in two-dimensions) to grow slightly faster than $t^{1/2}$. This is consistent with a radial growth rate faster than $t^{1/4}$ and is perhaps consistent with our theoretical prediction.
@@
In the numerical scheme, the initial sizes and positions of nuclei in a fully periodic domain are prescribed. While the positions remain fixed, the sizes of the nuclei evolve with time via diffusion of particles through the majority phase. These effective diffusive interactions between nuclei, responsible for Ostwald ripening, can be calculated analogously to the electrostatic energy of a collection of point charges [31, 33]. This perspective, while initially developed for passive systems, only relies on thermodynamic assumptions by employing an equilibrium Gibbs-Thomson relation as a boundary condition between each nucleus and the majority phase. Our derived Eqs. (6) are the analogue of the Gibbs-Thomson relation for an active system, allowing us to extend the perspective of Ref. [31] to active matter.
@@
For demonstration, we implement the active Voorhees-Glicksman theory to simulate the ripening of a periodic cell of active nuclei. Simulations of $250$ nuclei with random positions [50] and sizes were executed [33] with the nuclei separated by exceedingly large distances. It is therefore expected for the coarsening dynamics to be in agreement with LSW theory [11]. Noting again that the average radius is proportional to the critical radius, we expect the average radius across multiple activities to collapse to the same curve when scaled by the prefactor of the growth rate found in Eq. (15), ${K\equiv\left(\rho^{\rm sat}\upgamma_{\rm ost}d_{\rm hs}/\zeta U_{o}\Delta\rho^{2}\right)^{1/3}}$. The average nucleus size from these simulations scaled by $K$ are shown in Fig. 5. Our numerical simulations reveal that the average nucleus size is an excellent agreement with the predicted $\sim t^{1/3}$ scaling with curves for both bubbles and droplets collapsing for all activities. Plots of the distribution of nucleus sizes, which display evidence of evolving towards an LSW distribution [11], are provided in Section 4 of the SI [33].
@@
Figure 5: Average radius growth as a function of time for nuclei initialized with an LSW distribution [11] and average radius $750d_{\rm hs}$. Radius growth curves across different activities, as well as between bubbles and droplets, collapse onto one another when scaled by prefactor ${K\equiv\left(\rho^{\rm sat}\upgamma_{\rm ost}d_{\rm hs}/\zeta U_{o}\Delta\rho^{2}\right)^{1/3}}$.
@@
Although the Voorhees-Glicksman theory can be conveniently applied to late-stage active coarsening dynamics via Eqs. (6), it has significant limitations. In particular, the positions of the nuclei are fixed in time while in reality nuclei may diffuse, which may have significant impact on the inter-nuclei diffusion interactions and in extreme cases may lead to collision of nuclei (coalescence). More sophisticated coarsening theories [13] attempting to account for these effects, such as the work of Enomoto et al. [51] also only draw on thermodynamic arguments in the form of a Gibbs-Thomson relation. Using our work, such theories may then be extended to active matter so long as the non-equilibrium analogue to the Gibbs-Thomson relation employed in those works is established.

### v2

@@ II.2 Mechanics of a Critical Nucleus
Given a saturation density, one may use Eqs. (6) to solve for the unknowns $\rho^{\rm in}$ and $R_{C}$. We note that as $\rho^{\rm sat}$ approaches either binodal density, $R_{C}$ will diverge; Eq. (4a) and Eq. (5a) then simplify to the macroscopic phase coexistence criteria presented in Ref. [34]. As one approaches this limit and $R_{C}$ grows, $\mathcal{O}\left(R_{C}^{-2}\right)$ corrections to the coexistence criteria become negligible and we expect our theory to be most accurate. Solution of the coexistence criteria is complicated by the presence of $\upgamma_{\rm mech}$ and $\upgamma_{\rm ost}$ in Eqs. (6), which depend on the density profile, which in turn depends on $R_{C}$ and $\rho^{\rm in}$. Therefore, solution of Eqs. (6) requires self-consistently solving for the density profile using a procedure detailed in Section 1 of the SM. In Sec. II.3, we apply our mechanical theory to a canonical model system for active matter, active Brownian particles.
@@ II.3 Nucleation of Motility-Induced Phases
In three dimensions, the form of $\mathcal{P}(\rho)$ and $\{\kappa_{i}(\rho)\}$ for ABPs are given by [34, 26]:
$\mathcal{P}(\rho)$ $=p_{C}(\rho)+\frac{\rho\zeta U_{o}\ell_{o}\overline{U}(\rho)}{6},$ (8a)
$\kappa_{1}(\rho)$ $=\frac{\ell_{o}^{2}}{20}\left(\overline{U}(\rho)\right)^{2}\frac{\partial p_{C}}{\partial\rho},$ (8b)
$\kappa_{3}(\rho)$ $=\frac{\ell_{o}^{2}}{20}\overline{U}(\rho)\frac{\partial}{\partial\rho}\left[\overline{U}(\rho)\frac{\partial p_{C}}{\partial\rho}\right],$ (8c)
@@
$\mathcal{K}$ $=0,$ (8e)
where $p_{C}(\rho)$ is the conservative pressure arising from particle interactions and $\overline{U}(\rho)$ is the dimensionless effective active speed. In Eqs. (8) we have ignored the gradient contributions of the interaction potential to $\bm{\Sigma}$, which is justified when $\ell_{o}/d_{\rm hs}\gg 1$ for purely repulsive interactions [30, 34, 47]. Given our expressions for $\{\kappa_{i}(\rho)\}$, one can solve for the generalized Maxwell construction pseudovariable $\mathcal{E}\sim p_{C}(\rho)$ [34]. We also note that because $\mathcal{K}=0$, Eq. (4a) reduces to the familiar Laplace pressure difference (i.e. no non-local corrections) and the negative sign of $\upgamma_{\rm mech}$ [21] implies that the pressure inside of a droplet is lower than the pressure outside of the droplet [30]. The phase diagram, dependent on $\ell_{o}/d_{\rm hs}$ and volume fraction $\phi\equiv\rho\pi d_{\rm hs}^{3}/6$, has been well established computationally [5] and theoretically [34] with the latter predicting a critical point at $\phi\approx 0.49$ and $\ell_{o}\approx 16.2d_{\rm hs}$
Figure 2: Critical radius and relative change in nucleus density (${\nu^{\rm in}\equiv|\rho^{\rm in}-\rho^{\rm gas,\rm liq}|/\rho^{\rm gas,\rm liq}}$) as a function of supersaturation (${\nu^{\rm sat}\equiv|\rho^{\rm sat}-\rho^{\rm gas,\rm liq}|/\rho^{\rm gas,\rm liq}}$), as calculated from Eq. (6) using the numerical procedure outlined in Section 1.3 of the SM. Each data point requires self-consistent solution of the corresponding full spatial profile, e.g. those shown Fig. 1. Linear dependence of $\nu^{\rm in}$ on $\nu^{\rm sat}$ and inverse dependence of $R_{C}/d_{\rm hs}$ on $\nu^{\rm sat}$ implies a Gibbs-Thomson relation with $\nu^{\rm{in}}\propto 1/R$
@@
Figure 2 displays a striking difference between critical bubbles and droplets. At all examined values of activity, critical droplets of active Brownian particles are over an order of magnitude larger than critical bubbles under similar saturation conditions. This difference between bubbles and droplets widens as activity is increased and the phases become increasingly distinct. To contextualize this order of magnitude difference, we calculate the critical nucleus size and density for an equilibrium van der Waals fluid. While the size of the critical droplets are indeed larger than the size of critical bubbles in passive fluids, they remain within an order of magnitude of each other at similar saturation conditions even far from the critical point. In equilibrium, this difference is rooted in the asymmetry of the equations of state outside of the binodal densities: the vapor bubbles are more compressible than the liquid droplets. This asymmetry results in larger Laplace pressure differences for bubbles in comparison to droplets under similar saturation conditions, leading to the critical radius discrepancy. While this asymmetry in the compressibility is also present for active systems, the coexistence criteria contain uniquely nonequilibrium contributions, including a significant discrepancy between the vapor and bubble Ostwald tensions. At equilibrium these two quantities are identical, but can differ by an order of magnitude in active systems as shown in Fig. 3.
For a more physical understanding of the discrepancy in critical nucleus size between bubbles and droplets, we draw on the kinetic arguments independently proposed by Redner et al. [2] and Lee [22]. Establishing a flux-free critical nucleus with a density (and therefore interaction pressure) distinct from its surroundings requires the generation of an interfacial region of particles with a polarization pointing into the dense phase. Crucially, the rate at which particles with this orientation arrive to the interface must equal the rate at which they reorient and escape in order to satisfy the flux-free condition. The arrival rate is controlled by gas phase active speed and density while the escape rate is controlled by the minimum angle a particle must reorient in order to escape, the so-called “horizon angle [2].” For the same $R_{C}$, the rate of escape will depend greatly on whether the nucleus is the liquid phase (small horizon angle) or the gas phase (large horizon angle). As a result, for a similar supersaturation, the droplet must have a larger critical radius in order to satisfy the flux-free condition.
@@ II.4 Langevin Dynamics of a Nucleus
where $\rho$ now is a stochastic density field (while in the previous sections it was an ensemble average), $\bm{\Sigma}$ is defined identically as the dynamic stress tensor of Sec. II.3 and $\bm{\eta^{\rm act}}$ is an athermal stochastic contribution to the flux with zero mean and a variance of:
@@
We recently established a surface-area minimizing principle for liquid-gas interfaces of ABPs resulting from a positive capillary-wave tension, $\upgamma_{\rm cw}$, which resists fluctuations to the area-minimizing interfacial shape. The strength of the fluctuations were found to be set by the athermal active energy scale $k_{B}T^{\rm act}$ [26]. The ratio of the capillary tension to this energy scale – the active interfacial stiffness – controls the amplitude of long-wavelength interfacial fluctuations. A fluctuating nucleus can be well-approximated as a sphere when the average radius $R$ satisfies $(R^{2}\upgamma_{\rm cw}/k_{B}T^{\rm act})^{1/2}\gg 1$. Under this condition, the stochastic density field dynamics can be connected to the radial growth dynamics of the nucleus through the following radially symmetric ansatz [49, 4]:
@@
The radial growth dynamics take the form of a one-dimensional overdamped Langevin equation containing three distinct forces: a drag force with drag coefficient $\zeta_{\rm eff}$, a stochastic force with an athermal temperature scale of $k_{B}T^{\rm act}$ [50], and a deterministic force $F^{D}(R)$ proportional to the Ostwald tension. This deterministic force vanishes when $R=R_{C}$ where the density outside nucleus is identical to the far away saturation density $\rho^{\rm out}=\rho^{\rm sat}$. Crucially, the noise statistics satisfy the fluctuation-dissipation theorem (FDT) despite the underlying microscopic dynamics [Eqs. (7)] violating FDT. We can plainly see that $\upgamma_{\rm ost}$ is the quantity controlling the dynamics of nucleation and ripening. As first argued by Thjung et al. [23] a positive $\upgamma_{\rm ost}$ implies $R_{C}$ is an unstable fixed point in which nuclei with $R>R_{C}$ tend to grow while nuclei with $R<R_{C}$ tend to shrink. A priori, there is no reason to assume $\upgamma_{\rm ost}$ is generally positive. In the case of a negative $\upgamma_{\rm ost}$, $R_{C}$ becomes a stable fixed point. In such a scenario, one may expect phenomonology ranging from microphase separation to a “bubbly liquid” phase coexisting with a gas phase [23]. However, for ABPs we do not find a negative $\upgamma_{\rm ost}$ [see Fig. 4] – consistent with existing simulation work on three-dimensional ABPs [51, 5, 52, 26]. In two dimensions, some simulations of ABPs have reported phenomenology consistent with bubbly liquids [51, 53]. Whether these bubbly phases are a result of a negative $\upgamma_{\rm ost}$ or other effects [54] remains an open question.
@@
with $F^{D}(R)=-\partial U/\partial R$. In defining this potential, we have treated $\upgamma_{\rm ost}$ as independent of $R$ (just as in Ref. [4]), an approximation that appears to be well justified while $R\sim R_{C}$. This potential may be rearranged into a form familiar to equilibrium intuition:
@@
At equilibrium, $\mathcal{V}$ simplifies to the expected $\mu(\rho^{\rm sat})\Delta\rho-\Delta f$. From Eq. (13), one can clearly appreciate that the barrier contains two contributions: a positive term which depends solely on interfacial properties scaling as $R^{2}$ and a negative term which depends solely on bulk properties scaling as $R^{3}$. Thus the familiar picture of homogeneous nucleation where interfacial effects compete with bulk effects in order to define a critical radius and nucleation barrier is recovered but now contains generalized volumetric driving and interfacial penalty terms.
@@
The recovery of an effective detailed balance with respect to the size of a sub-critical nucleus directly implies that the probability of observing a nucleus of size $R<R_{C}$ is proportional to ${\text{exp}\left[-U(R)/k_{B}T^{\rm act}\right]}$ [4]. We note that the potential defined here relied on the assumption that the density profiles and quantities such as $\upgamma_{\rm ost}$ are constant with $R$ an approximation that clearly breaks down when $R<<R_{C}$. Even in equilibrium, the nucleation potential formulated using bulk thermodynamic quantities (e.g  the so-called ”capillary approximation”) is often used for all $R<R_{C}$ despite likely being beyond the range of applicability of bulk thermodynamic quantities [12, 55, 56]. Nevertheless, use of this potential for all $R$ allows one to straightforwardly extend the arguments of Becker-Döring theory [9, 12, 15] to extract the following kinetic expression for the rate at which an active droplet nucleates per unit volume, as outlined in Appendix A:
@@
The characteristic timescale to observe a nucleation event is given by $\tau_{\rm nuc}=(V\mathcal{R})^{-1}$ where $V$ is the total volume of the system. Thus the work of Redner et al. [2], which inferred the existence of a nucleation barrier by observing clustering statistics consistent with Becker-Döring theory from particle simulation, is well-justified on theoretical grounds, although our calculated barrier is for three dimensions while the simulations of Ref. [2] are in two. We note that the rate [Eq. (14)] is applicable to the nucleation of active droplets but not to active bubbles – the physical interpretation of the “clusters” posited in the derivation of Eq. (14) becomes unclear [12] in this context.
@@ II.5 Active Coarsening Dynamics
While Eqs. (6) and Eqs. (11) were motivated to describe the earliest stages of nucleation, i.e. the structure and dynamics of a single nucleus forming in a supersaturated solution, they may also describe late-stage coarsening dynamics long after all nucleation events have occurred. At this stage, the system will contain several nuclei of different sizes. The coarsening of these domains can proceed through a combination of several different mechanisms, including evaporation-condensation (Ostwald ripening) and coalescence. At sufficiently low $\nu^{\rm sat}$, the nuclei will be separated by large enough distances that collisions between them are exceedingly rare and coarsening will be dominated by Ostwald ripening. In such a scenario, Eqs. (11) can describe the dynamics of each nucleus such that they are only coupled via the density of the majority phase, $\rho^{\rm sat}$, which now evolves with time. The finite-size coexistence criteria presented in this work imply minor differences in density of the majority phase just outside nuclei of different sizes, inversely proportional to their radii. The majority phase therefore has an inhomogeneous density, driving transport of mass away from nuclei with enriched outside densities towards nuclei with depleted outside densities. For passive systems as well as ABPs [see Fig. 1], smaller droplets have enriched outside densities while smaller bubbles have depleted outside densities. Noting that transport of mass away from a nucleus would cause a droplet to shrink and a bubble to grow, this implies that large nuclei of minority phase will grow at the expense of smaller nuclei until the system reaches a final steady state with a single macroscopic domain. Nuclei compete to absorb (or inject) particles from the majority phase and therefore $R_{C}$, the radius below which nuclei shrink and above which nuclei grow, is now a function of time. While we do not have profiles for a particular system in which $\upgamma_{\rm ost}<0$, we suspect they would reveal smaller droplets having depleted outside densities and smaller bubbles having enriched outside densities. This would result in smaller nuclei growing at the expense of larger ones, consistent with stable radial dynamics of Eq. (11).
With the above physical picture in mind, we can determine the time-dependence of the post-nucleation critical radius in a similar manner as the Lifshitz-Slyozov-Wagner (LSW) theory [11]. In the limit of a vanishing phase fraction of the minority phase, the nuclei can be approximated as independent with each satisfying the Langevin equation shown in Eq. (11). The Fokker-Planck equation consistent with Eq. (11) can be straightforwardly found and its long-time asymptotic solution (while discarding terms of higher order than $R^{-3}$) can be determined analogously to LSW theory, as detailed in Appendix B. In doing so, we find that the critical radius evolves in time according to:
@@
where $\mathcal{C}$ is a dimensionless constant and $t_{o}$ is an arbitrary time origin. We note that this long-time asymptotic solution also implies that the average radius is always within a dimensionless multiplicative constant of $R_{C}$ [11, 14]. Equation (15) demonstrates that the $t^{1/3}$ growth law predicted by LSW theory should generically hold out of equilibrium while $\upgamma_{\rm ost}>0$. This $1/3$ power law has been measured experimentally in active systems [6, 7] with disparate microscopic details, suggesting that this growth law may in fact be robust. We note that the simulations of Redner et al. [8] measured the average number of particles per droplet (in two-dimensions) to grow slightly faster than $t^{1/2}$ at late times. This is consistent with a radial growth rate faster than $t^{1/4}$ and is perhaps consistent with our theoretical prediction.
Curiously, Ref. [8] also measured a much faster coarsening rate at early times where nuclei are spaced closely. While in this work we do not consider coarsening driven by nucleus collision, it is well-established for passive systems to also result in a $t^{1/3}$ growth law [57]. A transition from fast coarsening when nuclei are closely spaced to slow coarsening when nuclei are distant is eerily reminiscent of coarsening in systems with hydrodynamic interactions, in which collisions of nuclei (and subsequent relaxation to spherical geometry) induce flows that cause further collisions [58, 59]. While standard ABP models do not consider hydrodynamic interactions, a detailed accounting of the polarization field at the early stages of coarsening may reveal non-diffusive contributions. Furthermore, if $\ell_{o}$ greatly exceeds the spacing between nuclei, particles can be ballistically exchanged between nuclei, an effect that is also not accounted for in our perspective. A systematic treatment of early-stage active coarsening would help bridge the gap in our understanding between nucleation and late-stage coarsening and is the subject of future work.
@@
In the numerical scheme, the initial sizes and positions of nuclei in a fully periodic domain are prescribed. While the positions remain fixed, the sizes of the nuclei evolve with time via diffusion of particles through the majority phase. Additionally, the majority phase will relax much faster than the individual nuclei grow or shrink. We therefore consider the diffusive dynamics of only the majority phase (i.e., excluding the regions occupied by the nuclei) with the nuclei simply acting as particle generation terms (sources or sinks of particles). Considering only the majority phase dynamics implies that large density-gradient interfaces are ignored; we therefore discard gradient terms from $\bm{\Sigma}$ and arrive at the quasi-stationary expression for the majority phase’s density field (see Appendix C for further details):
$0=\bm{\nabla}^{2}\mathcal{P}(\rho)+\zeta\sum_{i=1}^{N_{\rm nuc}}S_{i}(\mathbf{r}),$ (16)
where $N_{\rm nuc}$ is the number of nuclei and $S_{i}(\mathbf{r})$ is a function quantifying its injection/absorption of particles into the majority phase. Using Eq. (16) to solve for the pressure field is analogous to solving for the electrostatic potential resulting from $N_{\rm nuc}$ charge densities. Approximating each nucleus as spherically symmetric allows one to treat the “charge densities” as “point charges”, i.e. $S_{i}(\mathbf{r})\sim\delta(\mathbf{r}-\mathbf{r}_{i})$. Then the effective diffusive interactions between nuclei, responsible for Ostwald ripening, can be calculated analogously to the electrostatic energy of a collection of point charges [31].
@@
This perspective, while initially developed for passive systems, only relies on thermodynamic assumptions by employing an equilibrium Gibbs-Thomson relation as a boundary condition between each nucleus and the majority phase, allowing for the determination of the magnitude of each $S_{i}$. Our derived Eqs. (6) are the analogue of the Gibbs-Thomson relation for an active system, allowing us to extend the perspective of Ref. [31] to active matter.
@@
For demonstration, we implement the active Voorhees-Glicksman theory to simulate the ripening of a periodic cell of active nuclei. Simulations of $250$ nuclei with random positions [60] and sizes were executed with the nuclei separated by exceedingly large distances. It is therefore expected for the coarsening dynamics to be in agreement with LSW theory [11]. Noting again that the average radius is proportional to the critical radius, we expect the average radius across multiple activities to collapse to the same curve when scaled by the prefactor of the growth rate found in Eq. (15), ${K\equiv\left(\rho^{\rm sat}\upgamma_{\rm ost}d_{\rm hs}/\zeta U_{o}\Delta\rho^{2}\right)^{1/3}}$. The average nucleus size from these simulations scaled by $K$ are shown in Fig. 5. Our numerical simulations reveal that the average nucleus size is an excellent agreement with the predicted $\sim t^{1/3}$ scaling with curves for both bubbles and droplets collapsing for all activities. Plots of the distribution of nucleus sizes, which display evidence of evolving towards an LSW distribution [11], are provided in Section 4 of the SM.
@@
Figure 5: Average radius growth as a function of time for nuclei initialized with an LSW distribution and average radius $750d_{\rm hs}$. Radius growth curves across different activities, as well as between bubbles and droplets, collapse onto one another when scaled by prefactor ${K\equiv\left(\rho^{\rm sat}\upgamma_{\rm ost}d_{\rm hs}/\zeta U_{o}\Delta\rho^{2}\right)^{1/3}}$.
@@
Although the Voorhees-Glicksman theory can be conveniently applied to late-stage active coarsening dynamics via Eqs. (6), it has significant limitations. In particular, the positions of the nuclei are fixed in time while in reality nuclei may diffuse, which may have significant impact on the inter-nuclei diffusion interactions and in extreme cases may lead to collision of nuclei (coalescence). More sophisticated coarsening theories [13] attempting to account for these effects, such as the work of Enomoto et al. [61] also only draw on thermodynamic arguments in the form of a Gibbs-Thomson relation. Our work establishes a Gibbs-Thomson relation for single-component phase separation purely from mechanics, without invoking equilibrium assumptions. Then established perspectives on classical coarsening may be extended to nonequilibrium systems as long as a nonequilibrium Gibbs-Thomson relation, such as that found in this work, is used.